i=v/z given v =-1+7.4j gven z = 3.1+5.1j
1 What is the modulus of the current
2 what is the argument of the current in radians

Answer by rfadrogane(214)

Before we get the modulus, lets convert first the given from rectangular form to polar form:
Use Pythagorean relation to solve for each magnitude of z & v.
v = [(-1)^2 +(7.4)^2]^(1/2) = 7.46
for the angle = arctan (-7.4/1) = 97.69 degrees
v = 7.46 angle 97.69 deg.
z = [(3.1)^2 +(5.1)^2]^(1/2) = 5.97
for the angle = arctan (5.1/3.1) = 58.71 degrees
z = 5.97 angle 58.71 deg.
and, finally:
i = v/z
= (7.46 angle 97.69 deg.)/(5.97 angle 58.71 deg.)
i = 1.25 angle 38.98 deg.
therefore;
a. 1.25 Amp.
b. 38.98 degrees